Sunday 3 October 1999 8.00am EST
|During the last 24 hours Chandra passed through the 7th eclipse of the season, exited the radiation belts, conducted an ACIS CTI measurement and began a series of HRC calibration measurements of AR Lac. These measurements will continue throughout the next orbit. The HRC door was re-opened after radiation belt exit. The spacecraft systems continued to operate nominally.
The current mission schedule command load began executing at 10.30AM 10/2 and will continue until 2.30PM 10/8.
Return to Status Reports